OpenStudy (anonymous):

solve: square root(x+7) + square root(x) = 2

OpenStudy (anonymous):

\[\sqrt{x+7}+\sqrt{x}=2\]

myininaya (myininaya):

\[\sqrt{x+7}=2-\sqrt{x}\]

myininaya (myininaya):

square both sides

OpenStudy (anonymous):

i got x = -3/2 lol

myininaya (myininaya):

\[x+7=4-2(2\sqrt{x})+x\]

myininaya (myininaya):

\[7=4-4\sqrt{x}\]

myininaya (myininaya):

\[3=-4\sqrt{x}\]

myininaya (myininaya):

\[\frac{3}{-4}=\sqrt{x}\]

myininaya (myininaya):

\[x=\frac{9}{16}\]

OpenStudy (anonymous):

i was quite wrong..

myininaya (myininaya):

we need to check it though

myininaya (myininaya):

it might not even be a solution since we squared both sides

myininaya (myininaya):

i will leave the checking to you just plug in 9/16 and see if both sides are the same